Process time Inspection time Move time Wait time 15.0 minutes 2.0 minutes 6.4 minutes 36.6 minutes Required a. Compute manufacturing cycle time b. Compute manufacturing cycle efficiency Complete this question by entering your answers in the tabs below Required A Required B Compute manufacturing cycle time. (Round your answers to 1 decimal place.) Manufacturing cycle time Time 15.0 minutes 2.0 minutes 6.4 minutes 36.6 minutes rocess time Inspection time Move time ait time Manufacturing cycle time 60.0 minutes
Answer #1

Manufacturing cycle time = Process time + Inspection time + move time + Wait time

                                    = (15.0+2.0+6.4+36.6)m = 60.0 m

Manufacturing cycle efficiency = Value-added time (process time) / Manufacturing cycle time

                                   MCE = 15.0 m / 60.0 m = 0.25 m = 25%